OpenStudy (wade123):

Write the sum using summation notation, assuming the suggested pattern continues. -3 + 6 + 15 + 24 + ... + 132

OpenStudy (wade123):

my answer:

OpenStudy (anonymous):

580.5 right....?

OpenStudy (wade123):

what?

OpenStudy (anonymous):

a=-3 c.d d=15-6=9 l=a+(n-1)d 132=-3+(n-1)9 132+3=(n-1)9 n-1=135/9=15 n=15+1=16 so your answer is correct.